# Computing Large 2D/3D Convex Regions of Obstacle-Free Space for Robotic Planning

### v1

@@ III-B Proposed Algorithm
A sequence $\{{\mathcal{P}_{1},\mathcal{E}_{1},\dots,\mathcal{P}_{k},\mathcal{E}_{k},\dots}\}$ is generated by repeating the iteration. Since $\operatorname*{vol}({\mathcal{E}_{k}})$ is non-decreasing and bounded, the ellipsoid volume will converge to a finite value. Consequently, FIRI terminates when this lower bound for $\operatorname*{vol}({\mathcal{P}_{k}})$ cannot be sufficiently improved (line 1).
Algorithm 1 describes a skeleton of FIRI. In conclusion, Restrictive Inflation brings managibility to FIRI, and monotonically iterative updates allows FIRI to continuously optimize the lower bound of (3) to obtain a high quality output. It is worth emphasizing that the efficiency of FIRI relies strongly on the performance of solving two optimization programming (line 1 and 1). In subsequent sections, we propose novel subalgorithms that are practically efficient and reliable, exploiting the geometrical structure of the involved programming, which greatly benefits the computational efficiency of FIRI.
IV Solving Restrictive Halfspace Computation via SDQP
Combining the definition of the halfspace $\mathcal{H}({a_{i}})$ in (9), we formulate the halfspace calculation in Restrictive Inflation defined in line 1 of Algorithm 1 into
@@ IV Solving Restrictive Halfspace Computation via SDQP
Although such a inflated ball $\mathcal{B}_{i}$ maximization is nonconvex, we can obtain its equivalent convex QP formulation through its polar duality. Algorithm 1 keeps $a_{i}^{\mathrm{T}}a_{i}>0$, which ensures that the origin always lies within the interior of the the halfspace $\mathcal{H}({a_{i}})$. Thus we directly optimize the polar dual vector of $a_{i}$ by substituting $a_{i}=b/(b^{\mathrm{T}}b)$ and obtain an equivalent QP
@@
$~{}u^{\mathrm{T}}b\geq 1,~{}\forall u\in\bar{\mathcal{O}}_{i}.$ (16c)
Note that (IV) is a $L_{2}$-norm minimization problem after inverse affine map in Algorithm 1 (line 1 and line 1), whereas it takes the general form of a strictly convex QP before map, and its positive definite matrix of the quadratic objective function is determined by the ellipsoid matrix of current iteration. In other words, the first inner loop of Algorithm 1 implicitly transforms a strictly convex QP into an equivalent form of $L_{2}$-norm minimization.
To efficiently perform the restrictive halfspace computation, we propose a randomized method, SDQP, for strictly convex QPs with small dimension but massive constraints, which computes exact solutions. This method generalizes Seidel’s algorithm from Linear Programming (LP) to strictly convex QP, and enjoys linear expected complexity about the constraint number. Then we provide in detail the flow of SDQP: 1) converting a strictly convex QP into $L_{2}$-norm minimization, 2) solving the minimization analytically in linear complexity.
We consider the following strictly convex QP,
$\min_{x\in\mathbb{R}^{n}}{\frac{1}{2}x^{\mathrm{T}}M_{\mathcal{Q}}x+c_{\mathcal{Q}}^{\mathrm{T}}x},~{}\mathrm{s.t.}~{}A_{\mathcal{Q}}x\leq b_{\mathcal{Q}},$ (17)
where $M_{\mathcal{Q}}\in\mathbb{R}^{n\times n}$ is positive definite, $c_{\mathcal{Q}}\in\mathbb{R}^{n}$, $A_{\mathcal{Q}}\in\mathbb{R}^{d\times n}$, and $b_{\mathcal{Q}}\in\mathbb{R}^{d}$. $d$ denotes the number of constraints, much larger than $n$. Thus the Cholesky factorization of $M_{\mathcal{Q}}$ exists and is unique, which we denote as $M_{\mathcal{Q}}=L_{\mathcal{Q}}L_{\mathcal{Q}}^{\mathrm{T}}$. We introduce a new variablel $y\in\mathbb{R}^{n}$ that satisfies
$y=L_{\mathcal{Q}}^{\mathrm{T}}x+L_{\mathcal{Q}}^{-1}c_{\mathcal{Q}}.$ (18)
Then the original QP (17) is equivalent to a $L_{2}$-norm minimization below,
$\min_{y\in\mathbb{R}^{n}}{\frac{1}{2}y^{\mathrm{T}}y},~{}\mathrm{s.t.}~{}Ey\leq f,$ (19)
where $E=A_{\mathcal{Q}}L_{\mathcal{Q}}^{-\mathrm{T}}$ and $f=A_{\mathcal{Q}}(L_{\mathcal{Q}}L_{\mathcal{Q}}^{\mathrm{T}})^{-1}c_{\mathcal{Q}}+b_{\mathcal{Q}}$. It is apparent that solving the minimization (19) can directly yield the solution to the associated QP (17), and this transformation is implicitly executed in Algorithm 1. Hereafter, we present a randomized algorithm with linear complexity for the low-dimensional $L_{2}$-norm minimization (19) in Algorithm 2.
Input: $\mathcal{H}_{E}$
@@
2       $y\leftarrow\mathbf{0}$ if $dim(\mathcal{H}_{E})==1$ then
3             $y\leftarrow\mathbf{OneDimMinNorm}(\mathcal{H}_{E})$
@@
5      ${}\mathcal{I}\leftarrow\{\}$ foreach $h\in\mathcal{H}_{E}~{}~{}\mathrm{in~{}a~{}random~{}order}$ do
@@
7                   $\{M,v,{\mathcal{H}_{E}}^{\prime}\}\leftarrow\mathbf{HouseholderProj}(\mathcal{I},h)$ $y^{\prime}\leftarrow\mathbf{LowDimMinNorm}({{\mathcal{H}_{E}}^{\prime}})$ $y\leftarrow My^{\prime}+v$
@@ 9            $\mathcal{I}\leftarrow\mathcal{I}\cup\{h\}$
13
Algorithm 2 LowDimMinNorm
As the input of Algorithm 2, we define $\mathcal{H}_{E}$ as the set of the hyperplanes corresponding to the constraints of the $L_{2}$-norm minimization (19). The central idea of Algorithm 2 is to start with $y=\mathbf{0}$ which is the solution of an unconstrained $L_{2}$-norm minimization (line 2), and then gradually check the constraint $h\in\mathcal{H}_{E}$ in random order (line 2), denoting $I$ as the set of the constraints that have already been checked, as illustrated in Fig. 2(a). We check whether the solution under the constraints $I$ violates the new constraint $h$. If it is not violated as shown in Fig. 2(b), next constraint will be checked. If it is violated as shown in Fig. 2(c), $h$ must be active for the $L_{2}$-norm minimization under constraints $\mathcal{I}\cup\{h\}$. That is, the solution must be on the constraint plane of $h$, thus minimization of this iteration can be written as
$\min_{y\in\mathbb{R}^{n}}~{}$ ${\frac{1}{2}y^{\mathrm{T}}y},$ (20a)
@@
where $E_{\mathcal{I}}$ and $f_{\mathcal{I}}$ represent the coefficients corresponding to the hyperplanes in $\mathcal{I}$, $E_{h}$ and $f_{h}$ represent the coefficients of $h$. Using the geometric structure of this problem, we transform the minimization into a subproblem of $n-1$ dimensional $L_{2}$-norm minimization with the same form as (19) as shown in Fig. 2(d), which will then be presented in detail. This allows for a recursive call of Algorithm 2 (line 2). When $n=1$, the problem (19) is equivalent to a trivial problem of computing the smallest absolute value in a interval, and the feasibility determination of the problem and solution can be calculated directly. We assume that the subproblem can be successfully solved, and thus the solution of this iteration can be calculated (line 2), then the next constraint can be checked.
Figure 2: Illustration of Algorithm 2 in 2-D. (a) indicates that both the inequality constraints $h_{1}$ and $h_{2}$ have been checked, which means $h_{1},h_{2}\in\mathcal{I}$. And $y$ is the solution of 2-D $L_{2}$-norm minimization under $\mathcal{I}$. (b) illustrates the case where $y$ does not violate the newly added constraint $h$. (c) shows the case where the original solution under $\mathcal{I}$ violates the new constraint $h$, thus we need to find a new solution $y$ in the constraint plane corresponding to $h$, implying equation constraint. The vector $v$ is normal to the constraint plane, and $H^{\mathrm{T}}e_{1}$, $H^{\mathrm{T}}e_{2}$ are a set of orthogonal basis of the 2-D space, where $H^{\mathrm{T}}e_{2}\bot v$. In (d), we establish a new coordinate for the 1-D subspace in the constraint plane with $v$ as the origin and $H^{\mathrm{T}}e_{2}$ as the orthogonal basis, transform the checked constraints to this coordinate as $h^{\prime}_{1},h^{\prime}_{2}$, and finally transforms (c) into a 1-D $L_{2}$-norm minimization with only inequality constraints.
An essential process of Algorithm 2 is the transformation of (IV) into a $n-1$ dimensional subproblem with the same form as problem (19) in line 2. The efficiency and numerical stability of this process have a great impact on the efficiency and stability of Algorithm 2. Since $L_{2}$-norm is invariant under the orthogonal transformation, we establish a new $n~{}-~{}1$ dimensional Cartesian coordinate system on the constraint plane (20c). As illustrated in Fig. 2(c), we take the point of $L_{2}$-norm minimization on the constraint plane as the origin of the coordinate system,
@@
$\dfrac{1}{2}y^{\mathrm{T}}y=\dfrac{1}{2}(y-v)^{\mathrm{T}}(y-v)+\dfrac{1}{2}v^{\mathrm{T}}v.$ (22)
which means that, with the constraint (20c), the $L_{2}$-norm minimization of $y$ is equivalent to the $L_{2}$-norm minimization of the $(y-v)$ which can be transformed into an $n-1$ dimensional vector in the newly established coordinate system.
Then we establish the new coordinate system by Householder reflection. $v$ is a normal vector of the constraint plane proportional to the geometric scale of the problem, which has intuitive numerical stability. Thus we use the normal vector $v$ to compute the Householder reflection. Taking the index of the element of $v$ with the largest absolute value as
$j=\mathrm{argmax}_{k\in\{1,...,n\}}\|v_{k}\|.$ (23)
Then the obtuse reflection vector $u$ that transforms $v$ to be parallel to $e_{j}$ is
$u=v+\operatorname{sgn}(v_{j})\|v\|e_{j},$ (24)
whose corresponding Householder matrix is
$H=\mathbf{I}_{n}-\dfrac{2uu^{\mathrm{T}}}{u^{\mathrm{T}}u},$ (25)
which is an orthogonal matrix. The use of the obtuse reflection vector $u$ corresponding to the dimension of the largest absolute value element of $v$ prevents the Householder transformation from being ill-conditioned due to a small reflection angle. This procedure is implicitly equivalent to a single-step operation of the Householder QR factorization. Easy to verify
$Hv=-\operatorname{sgn}(v_{j})\|v\|e_{j},$ (26)
which indicates that the set of all col vectors of the orthogonal $H^{\mathrm{T}}$ except the $j$-th one form an orthonormal basis for the orthogonal complement of $v$, as shown in Fig. 2(c). This process of calculating orthogonal basis has higher numerical stability than the Gram-Schmidt orthogonalization which may fall victim to the catastrophic cancelation problem [36]. Since $v$ is normal to constraint plane, we adopt this orthonormal basis to define the new coordinate system of the $n-1$ dimensional subspace on the constraint plane, and introduce the corresponding coordinates $y^{\prime}\in\mathbb{R}^{n-1}$. We denote $M\in{R}^{n\times(n-1)}$ as the matrix obtained by removing the $j$-th column of $H^{\mathrm{T}}$. Then the point $y$ in the origin coordinate system corresponding to $y^{\prime}$ is
@@
$\dfrac{1}{2}\|My^{\prime}\|^{2}=\dfrac{1}{2}{y^{\prime}}^{\mathrm{T}}{M}^{\mathrm{T}}{M}{y^{\prime}}=\dfrac{1}{2}{y^{\prime}}^{\mathrm{T}}{y^{\prime}},$ (28)
Eventually, for (IV) with a equality constraint, we construct its corresponding $n-1$ dimensional $L_{2}$-norm minimization with only linear inequality constraints on the constraint plane as
$\min\limits_{{y^{\prime}}\in\mathbb{R}^{n-1}}~{}$ $\dfrac{1}{2}{y^{\prime}}^{\mathrm{T}}{y^{\prime}},$ (29a)
@@
As Fig. 2(d) shows, the linear inequality constraints on ${y^{\prime}}$ can be obtained by substituting (27) into $E_{\mathcal{I}}y\leq f_{\mathcal{I}}$, and the set of their corresponding hyperplanes is denoted as ${\mathcal{H}_{E}}^{\prime}$ (line 2-2). Obviously, (IV) has the same structure as (19) and has a lower dimension, thus it can be solved by recursively calling Algorithm 2 (line 2).
Finally, we complete the extension of Seidel’s algorithm [9] to strictly positive definite QP with inequality constraints. There is a probability to trigger a recursive calls of the $n-1$ dimensional subproblem in each iteration. The idea of this randomized algorithm is that when the dimension $n$ of the problem is small but the constraint scale $d$ is huge, the probability of a newly added constraint becoming a active constraint is tiny, hence the probability of recursive calls is tiny as well. The expected complexity of this algorithm algorithm is $O(n!d)$, so that in the common dimensions $n\in\{2,3\}$, the expected complexity only increases linearly with the constraint scale [9]. In addition, since the algorithm preprocess the order of constraints by a random permutation, its actual performance does not depend on the input order and almost always achieves expected linear complexity. With the help of this randomized algorithm, the computational complexity of handling all obstacles in the constrained inflation of Algorithm 1 grows linearly only with the total number of vertices of the obstacles.

### v2

@@ III-B3 Manageability and Convergence of FIRI
Then a sequence $\{{\mathcal{P}_{1},\mathcal{E}_{1},\dots,\mathcal{P}_{k},\mathcal{E}_{k},\dots}\}$ is generated by repeating the iteration. Since $\operatorname*{vol}({\mathcal{E}_{k}})$ is non-decreasing and bounded, the ellipsoid volume will converge to a finite value. Consequently, FIRI terminates when this lower bound for $\operatorname*{vol}({\mathcal{P}_{k}})$ cannot be sufficiently improved (Line 1).
In conclusion, RsI brings manageability to FIRI, and the monotonically iterative updates allows FIRI to continuously optimize the lower bound of the polytope’s volume to obtain a high quality output. It is worth emphasizing that the efficiency of FIRI relies strongly on the performance of solving two optimization programming in Line 1 and 1. In subsequent sections, we propose novel subalgorithms that are practically efficient and reliable, exploiting the geometrical structure of the involved programming, which greatly benefits the computational efficiency of FIRI.
IV Solving Restrictive Halfspace Computation in RsI via SDMN
IV-A Reformulation of Restrictive Halfspace Computation
In this section, we focus on designing efficient method for the computation of halfspaces in RsI, defined in Line 1 of Algorithm 1. Combining the definition of the halfspace $\mathcal{H}({a_{i}})$ in (9), we formulate the halfspace calculation into
@@ IV-A Reformulation of Restrictive Halfspace Computation
Algorithm 1 keeps $a_{i}^{\mathrm{T}}a_{i}>0$, which ensures that the origin always lies within the interior of the the halfspace $\mathcal{H}({a_{i}})$. Thus, although such a maximization of the inflated ball $\mathcal{B}_{i}$ is nonconvex, we can obtain its equivalent minimum-norm formulation through its polar duality. Specifically, we directly optimize the polar dual vector of $a_{i}$ by substituting $a_{i}=b/(b^{\mathrm{T}}b)$ and obtain an equivalent $L_{2}$-norm minimization
@@
$~{}u^{\mathrm{T}}b\geq 1,~{}\forall u\in\bar{\mathcal{O}}_{i},$ (15c)
which has a low-dimension but multi-constraint nature.
IV-B Solution to Small-Dimensional Minimum-Norm with Massive Constraints
For efficiently solving the new formulation (15) of restrictive halfspace computation, we propose an analytical method for such Small Dimensional Minimum-Norm, called SDMN. This method generalizes Seidel’s randomized algorithm [13] from Linear Programming (LP) to minimum-norm, and enjoys complexity linear in the constraint number.
Without loss of generality, we consider the following general small-dimensional minimum-norm problem,
$\min_{y\in\mathbb{R}^{n}}{y^{\mathrm{T}}y},~{}\mathrm{s.t.}~{}Ey\leq f,$ (16)
where $E\in\mathbb{R}^{d\times n}$, $f\in\mathbb{R}^{d}$, $d$ denotes the number of constraints, which is much larger than $n$. Hereafter, we present a randomized algorithm with linear complexity for the small-dimensional minimum-norm (16) in Algorithm 2.
As shown in Algorithm 2, this is a recursive algorithm. In the following, we first provide an outline of the algorithm, then we give a detailed description of how to construct the recursive problem, and finally present a complexity analysis.
Notion:  constraints already checked $\mathcal{I}$,
            input and output of the recursive call ${{\mathcal{H}_{E}}^{\prime}},~{}y^{\prime}$
Input:   set of halfspace constraints $\mathcal{H}_{E}$
@@ IV-B Solution to Small-Dimensional Minimum-Norm with Massive Constraints
2       $y\leftarrow\mathbf{0}$
3       if $\mathrm{dim}(\mathcal{H}_{E})==1$ then
4             $y\leftarrow\textbf{OneDimMinNorm}(\mathcal{H}_{E})$
5             return $y$
@@ 4             $y\leftarrow\textbf{OneDimMinNorm}(\mathcal{H}_{E})$
7      ${}\mathcal{I}\leftarrow\{\}$
8       foreach $h\in\mathcal{H}_{E}~{}~{}\mathrm{in~{}a~{}random~{}order}$ do
@@ 7      ${}\mathcal{I}\leftarrow\{\}$
10                   /* Detail in Sec. IV-B2 */
11                   $\{M,v,{\mathcal{H}_{E}}^{\prime}\}\leftarrow\textbf{HouseholderProj}(\mathcal{I},h)$
12                   $y^{\prime}\leftarrow\textbf{SDMN}({{\mathcal{H}_{E}}^{\prime}})$
13                   $y\leftarrow My^{\prime}+v$
14
@@ 16            $\mathcal{I}\leftarrow\mathcal{I}\cup\{h\}$
21
Algorithm 2 SDMN
IV-B1 Algorithm Outline
We denote $\mathcal{H}_{E}$ as the set of the hyperplanes corresponding to the constraints of the $L_{2}$-norm minimization (16), and denote $\mathcal{I}$ as the set of the constraints that have already been checked (Line 2). Algorithm 2 starts with $y=\mathbf{0}$ which is the solution of an unconstrained $L_{2}$-norm minimization (Line 2), and then gradually checks the constraint $h\in\mathcal{H}_{E}$ in a random order (Line 2). We check whether the solution under the constraints $\mathcal{I}$ violates the new constraint $h$. If it is not violated, the next constraint will be checked. If it is violated, $h$ must be active. That is, the solution must be on the boundary of $h$, thus the minimization (16) under the constraints $\mathcal{I}\cup\{h\}$ can be written as
$\min_{y\in\mathbb{R}^{n}}~{}$ ${y^{\mathrm{T}}y},$ (17a)
@@ IV-B1 Algorithm Outline
where $E_{\mathcal{I}}$ and $f_{\mathcal{I}}$ represent the coefficients corresponding to the hyperplanes in $\mathcal{I}$, $E_{h}$ and $f_{h}$ represent the coefficients of $h$. Using the geometric structure of the problem (17), we transform it into a subproblem of $(n-1)$ dimensional $L_{2}$-norm minimization with the same form as (16) in Line 2, which is described in detail in Sec. IV-B2. This transformation allows for a recursive call of Algorithm 2 (Line 2). Additionally, when $n=1$, the problem (16) is equivalent to a problem of computing the smallest absolute value in a interval (Line 2), whose solution can be calculated trivially. We assume that the subproblem can be successfully solved, and thus the new solution under the constraints $\mathcal{I}\cup\{h\}$ can be calculated (Line 2). Subsequently, the next constraint can be checked until all constraints in the set $\mathcal{H}_{E}$ are examined, from which we obtain the result of Algorithm 2. To provide an intuitive perception, we present an example of the violation check in Fig. 3, where we use $y_{old}$ and $y_{new}$ to distinguish between solutions under $\mathcal{I}$ and $\mathcal{I}\cup\{h\}$.
Figure 3: Illustration of a specific instance of Algorithm 2 in 2-D for Sec. IV-B1 and Sec. IV-B2. For Sec. IV-B1: (a): Both the inequality constraints $h_{1}$ and $h_{2}$ have been checked, which means =$\mathcal{I}=\{h_{1},h_{2}\}$. $y_{old}$ is the solution of the 2-D $L_{2}$-norm minimization under $\mathcal{I}$. (a)$\Rightarrow$(b): When $y_{old}$ does not violate the newly added constraint $h$, $y_{new}=y_{old}$. (a)$\Rightarrow$(c): When $y_{old}$ violates the new constraint $h$, we need to find a new solution $y_{new}$ on the constraint plane corresponding to $h$, implying equation constraint (17). For Sec. IV-B2: The vector $v$ is normal to the constraint plane, and $H^{\mathrm{T}}e_{1}$, $H^{\mathrm{T}}e_{2}$ are a set of orthogonal basis of the 2-D space, where $H^{\mathrm{T}}e_{2}\bot v$. (d): We establish a new coordinate for the 1-D subspace on the constraint plane with $v$ as the origin and $H^{\mathrm{T}}e_{2}$ as the orthogonal basis, transform the checked constraints $h_{1},h_{2}$ to this coordinate as $h^{\prime}_{1},h^{\prime}_{2}$. Finally, we transforms (c) into a 1-D $L_{2}$-norm minimization with only inequality constraints (23).
IV-B2 Recursive Problem Construction
An essential process of Algorithm 2 is the recursive problem construction (Line 2-2), which transforms (17) that has an equality constraint into a $(n-1)$ dimensional subproblem with the same form as the original problem (16). The efficiency and numerical stability of this process have a great impact on Algorithm 2. Since $L_{2}$-norm is invariant under the orthogonal transformation, we establish a new $(n-1)$ dimensional Cartesian coordinate system on the constraint plane (17c) to implement the construction of the reduced dimensional recursive problem. We provide a specific example of this process in Fig. 3(c) and 3(d). As illustrated in Fig. 3(c), we take the point of $L_{2}$-norm minimization on the constraint plane as the origin of the coordinate system,
@@ IV-B2 Recursive Problem Construction
$y^{\mathrm{T}}y=(y-v)^{\mathrm{T}}(y-v)+v^{\mathrm{T}}v,$ (19)
which means that, with the constraint (17c), the $L_{2}$-norm minimization of $y$ is equivalent to the $L_{2}$-norm minimization of $(y-v)$ which can be transformed into an $(n-1)$ dimensional vector in the newly established coordinate system.
We construct the new coordinate system by finding an orthogonal matrix $H\in\mathbb{R}^{n\times n}$ that satisfies the requirement
$Hv\propto e_{j},$ (20)
where $e_{j}\in\mathbb{R}^{n}$ is an unit vector whose $j$-th element is 1. We provide the details of the choice of $j$ and the computation of $H$ at the end of Sec. IV-B2. The requirement (20) indicates that the set of all column vectors of the orthogonal $H^{\mathrm{T}}$ except the $j$-th one form an orthonormal basis for the orthogonal complement of $v$, as shown in Fig. 3(c). Since $v$ is normal to the constraint plane, we adopt this orthonormal basis to define the new coordinate system of the $(n-1)$ dimensional subspace on the constraint plane, and introduce the corresponding coordinates $y^{\prime}\in\mathbb{R}^{n-1}$. We denote $M\in\mathbb{R}^{n\times(n-1)}$ as the matrix obtained by removing the $j$-th column of $H^{\mathrm{T}}$. Then $y$ in the origin coordinate system corresponding to $y^{\prime}$ can be computed as
@@
$y^{\mathrm{T}}y-v^{\mathrm{T}}v={y^{\prime}}^{\mathrm{T}}{M}^{\mathrm{T}}{M}{y^{\prime}}={y^{\prime}}^{\mathrm{T}}{y^{\prime}}.$ (22)
Eventually, for (17) with an equality constraint, we construct its corresponding $(n-1)$ dimensional $L_{2}$-norm minimization with only inequality constraints on the constraint plane as
$\min\limits_{{y^{\prime}}\in\mathbb{R}^{n-1}}~{}$ ${y^{\prime}}^{\mathrm{T}}{y^{\prime}},$ (23a)
@@
As Fig. 3(d) shows, the linear inequality constraints on ${y^{\prime}}$ can be obtained by substituting (21) into the inequality constraint (17b) of the original problem, and the set of their corresponding hyperplanes is denoted as ${\mathcal{H}_{E}}^{\prime}$ (Line 2-2). Obviously, (23) has the same structure as (16) and has a lower dimension, thus it can be solved by recursively calling Algorithm 2 (Line 2), which gives rise to the formation of a recursive algorithmic structure.
Here we provide the details of $j$ and $H$ in (20). We obtain the orthogonal matrix $H$ by Householder reflection [43]. $v$ is a normal vector of the constraint plane proportional to the geometric scale of the problem, which has intuitive numerical stability. Thus, we use the normal vector $v$ to compute the Householder reflection. First, we set $j$ as the index of the element of $v$ with the largest absolute value as
$j=\mathrm{argmax}_{k\in\{1,...,n\}}\|v^{\mathrm{T}}e_{k}\|.$ (24)
Then the reflection vector $u$ that transforms $v$ to be parallel to $e_{j}$ can be calculated by
$u=v+\operatorname{sgn}(v_{j})\|v\|e_{j},$ (25)
whose corresponding Householder matrix can be calculated by
$H=\mathbf{I}_{n}-\dfrac{2uu^{\mathrm{T}}}{u^{\mathrm{T}}u},$ (26)
which is an orthogonal matrix and satisfies the aforementioned requirement (20). The use of the obtuse reflection vector $u$ corresponding to the opposite direction of the largest absolute value element of $v$ prevents the Householder transformation from being ill-conditioned due to a small reflection angle. This procedure is implicitly equivalent to a single-step operation of the Householder QR factorization, which has higher numerical stability than the Gram-Schmidt orthogonalization which may fall victim to the catastrophic cancelation problem [44].
IV-B3 Complexity Analysis
To conclude this section, we give the complexity analysis of the randomized Algorithm 2. There is a probability to trigger a recursive call of the $(n-1)$ dimensional subproblem in each iteration. The idea of this randomized algorithm is that when the dimension $n$ of the problem is small but the constraint scale $d$ is huge, the probability of a newly added constraint becoming an active constraint is tiny, thus the probability of recursive calls is tiny as well. The expected complexity of this algorithm is $O(n!d)$ [13], so that in the common dimensions $n\in\{2,3\}$, the expected complexity only increases linearly with the constraint scale. In addition, since the algorithm preprocess the order of constraints by a random permutation, its actual performance does not depend on the input order and almost always achieves expected linear complexity. With the help of this randomized algorithm, the complexity of handling all obstacles in the RsI of Algorithm 1 grows linearly only with the total number of vertices of the obstacles and the seed.